OpenStudy (anonymous):

1. Solve the equation for y. 5x – 7y = 10 (Points : 1) A. y=-7/5x-7/10 B. y=5/7x+10/5 C. 5/7x-10/7 D. y=-5x-7

OpenStudy (ivettef365):

to isolate y first step as follows; 5x - 7y - 10 -7y = -5x -10 (change the 5x to this side as negative)

OpenStudy (anonymous):

why are you changing it to a negative?

OpenStudy (ivettef365):

then divide by -7 and you have -7y = -5x + 10 --- --- --- -7 -7 -7 then you havd: y = 5/7x -10/7

OpenStudy (anonymous):

thank you:D